1a. A weak monoprotic base has a Kb = 1.67 × 10−10. Calculate the Ka for...

1a.

A weak monoprotic base has a Kb = 1.67 × 10−10. Calculate the Ka for the conjugate acid of this base.

Report your answer to THREE significant figures. You may (but are not required to) enter your answer in scientific notation; for example, if your answer is 2.44 × 10-2, you would enter 2.44E-2. The E must be uppercase.

1b.

A weak monoprotic base has a Kb = 3.38 × 10−3. Calculate the Ka for the conjugate acid of this base.

Report your answer to THREE significant figures. You may (but are not required to) enter your answer in scientific notation; for example, if your answer is 2.44 × 10-2, you would enter 2.44E-2. The E must be uppercase.

1c.

A weak monoprotic acid has a Ka = 4.28 × 10−4. Calculate the Kb for the conjugate base of this acid.

Report your answer to THREE significant figures. You may (but are not required to) enter your answer in scientific notation; for example, if your answer is 2.44 × 10-2, you would enter 2.44E-2. The E must be uppercase.

1d.

A weak monoprotic acid has a Ka = 2.19 × 10−7. Calculate the Kb for the conjugate base of this acid.

Report your answer to THREE significant figures. You may (but are not required to) enter your answer in scientific notation; for example, if your answer is 2.44 × 10-2, you would enter 2.44E-2. The E must be uppercase.
Answer #1

la. Ko = 1.6781010 Chiven) As we know kaxk6 = Kw Kaxko = 1x10-14 =0.598 104 Ka=1810-14 1.678 0 10 = 5.98x 10 or 5.98 €-5 - K6